OpenStudy (anonymous):

Analytic Geometry Find the angle between the plane 2x-3y+z-4=0 and the line 3x+y+2z-1=0 , x+7y-2z+5=0 ??

OpenStudy (experimentx):

for plane ... find the direction of normal (or Direction Cosines) for two lines ... solve them to find parametric eqn or reduce them to (x-a)/l = (y-b)/m = (z-c)/n where l,m,n are direction cosines ... since you have direction cosines of both ... you can find the angle between them